### v1

# Probing Cosmic Curvature with Fast Radio Bursts and DESI DR2

Abstract
The spatial curvature of the Universe remains a central question in modern cosmology. In this work, we explore the potential of localized Fast Radio Bursts (FRBs) as a novel tool to constrain the cosmic curvature parameter $\Omega_{k}$ in a cosmological model-independent way. Using a sample of 80 FRBs with known redshifts and dispersion measures, we reconstruct the Hubble parameter $H(z)$ via artificial neural networks, which is then used to obtain angular diameter distances $D_{A}(z)$ through two complementary approaches. First, we derive the comoving distance $D_{C}(z)$ and $D_{A}(z)$ from FRBs without assuming a fiducial cosmological model. Then, the $H(z)$ reconstruction is combined with Baryon Acoustic Oscillation (BAO) measurements to infer $D_{A}(z)$. By comparing the FRB-derived and BAO+FRB-derived $D_{A}(z)$ values, we extract constraints on $\Omega_{k}$ under the Friedmann–Lemaître–Robertson–Walker metric. Our results show consistency with a spatially flat Universe within $1\sigma$ uncertainties, although a mild preference for negative curvature is observed across both covariance-based and Gaussian analyses. This study highlights the growing relevance of FRBs in precision cosmology and demonstrates their synergy with BAO data as a powerful probe of the large-scale geometry of the Universe.
Keywords: Fast Radio Bursts, Artificial Neural Networks, Cosmic Distances, Curvature of Universe.
I Introduction
Determining the spatial geometry of the Universe through the curvature parameter $\Omega_{k}$ is a fundamental goal in cosmology. It directly impacts the measurement of distances, as well as our understanding of spacetime, the global topology of the Universe, its ultimate fate, and the physics of the early times, for instance.
The current constraints on $\Omega_{k}$ are largely consistent with flatness, although uncertainties remain. For instance, combining Planck Cosmic Microwave Background (CMB) lensing data with low-redshift Baryon Acoustic Oscillations (BAO) measurements yields $\Omega_{k}=0.0007\pm 0.0019$ Aghanim et al. (2020). However, other analyses challenge this result. Notably, Di Valentino et al. (2020) reported a preference for a mildly closed Universe, with $\Omega_{k}=-0.044^{+0.018}_{-0.015}$, based solely on CMB temperature and polarization data from Planck satellite. Another study found a preference for a closed Universe at the $2\sigma$ confidence level, measuring $\Omega_{k}=-0.089^{+0.049}_{-0.046}$ when applying Effective Field Theories of Large Scale Structure (EFTofLSS) to BAO data Glanville et al. (2022). On the other hand, Wu and Zhang (2024) reported $\Omega_{k}=0.108\pm 0.056$, which is in $2.6\sigma$ tension with the CMB lensing data result. To obtain this estimate, they used a combination of datasets, including BAO, Cosmic Chronometers (CC), and Strong Gravitational Lensing (SGL).
However, it is important to recognize that these constraints are inherently cosmological model-dependent, relying on assumptions within the non-flat $\Lambda$CDM framework, which suffers from the well-known intrinsic degeneracy between cosmic curvature and the Hubble constant. Moreover, as with the Hubble tension, the apparent disagreement in curvature estimates raises the question of whether systematic biases or new physics are at play. In this context, exploring methods that provide a complementary and cosmological model-independent perspective is particularly valuable. Developing curvature measurements that are independent of the cosmological model, especially those that do not rely solely on CMB-based assumptions, offers a promising pathway to gaining a better understanding of these issues.
Recent studies Gong et al. (2024); Liu et al. (2025), which provide a methodological basis for the present work, have proposed model-independent approaches by combining BAO and Hubble parameter data. Although their methodologies are broadly similar, the former utilizes only BOSS and eBOSS BAO data and omits the covariance matrix, whereas the latter incorporates full covariance information and includes additional measurements from DESI DR1. Both studies find results consistent with flat spatial curvature within the reported uncertainties.
Extending these methodologies, we adopt a comparable framework to systematically assess the impact of including versus omitting the covariance matrix on cosmic curvature estimates. Additionally, we incorporate FRBs as novel and independent tracers to reconstruct the Hubble parameter $H(z)$, offering a complementary perspective to traditional probes such as supernovae or cosmic chronometers. This inclusion introduces a new observational channel based on dispersion measures, thereby enhancing the robustness and versatility of model-independent curvature constraints. We further combine this with the latest BAO data from BOSS, eBOSS, and DESI DR2 Karim et al. (2025).
FRBs have emerged as a promising new probe in cosmology. These events are characterized by millisecond-duration and highly luminous pulses in the radio frequency spectrum. First discovered in 2007 by Lorimer et al. Lorimer et al. (2007), FRBs exhibit large dispersion measures (DMs), that significantly exceed the expected contribution from the Milky Way. This excess provides strong evidence for their extragalactic, and potentially cosmological, origin Petroff et al. (2019). Since that initial detection, hundreds of FRBs have been observed, and a subset has been precisely localized to host galaxies with measured redshifts Zhou et al. (2022).
Beyond their intriguing astrophysical nature, FRBs have proven to be valuable probes for cosmology. In particular, localized FRBs, those with measured redshifts and host galaxy associated, provide unique observational access to the otherwise elusive ionized content of the intergalactic medium (IGM). Over the past few years, multiple studies have employed these events to estimate key cosmological quantities, such as the Hubble constant $H_{0}$ Macquart et al. (2020); Wu et al. (2022); Zhao et al. (2022); Kalita et al. (2025); Hagstotz et al. (2022); Fortunato et al. (2025), cosmographic parameters Fortunato et al. (2023), the cosmic proper distance Yu and Wang (2017), the baryon fraction in the IGM Li et al. (2020); Yang et al. (2022); Lemos et al. (2022), constrain dark energy equation of state Zhou et al. (2014); Gao et al. (2014), trace the magnetic fields in the IGM Akahori et al. (2016), probe the equivalence principle Wei et al. (2015); Tingay and Kaplan (2016); Nusser (2016), the rest mass of photons Wu et al. (2016); Shao and Zhang (2017); Lin et al. (2023), to cite some.
In this work, we implement a method to determine the cosmic curvature by leveraging the synergy between FRBs and BAO measurements, with FRBs serving as the core observable. By comparing two estimates of the angular diameter distance $D_{A}(z)$, independently reconstructed from FRBs alone and from FRBs+BAO data sets, we infer $\Omega_{k}$ without relying on a specific cosmological model. This is made possible by using the FRB dispersion measure to reconstruct $H(z)$ through an artificial neural network algorithm, as demonstrated in Fortunato et al. (2025) and to trace the comoving distance $D_{C}$ via the integrated electron density along the line of sight.
We assume the validity of the Cosmological Principle, adopting the premise that the Universe is homogeneous and isotropic on large scales. Consequently, the background spacetime is described by the Friedmann–Lemaître–Robertson–Walker (FLRW) metric, with the curvature parameter $\Omega_{k}$ determining the geometry of the Universe: flat ($\Omega_{k}=0$), open ($\Omega_{k}>0$), or closed ($\Omega_{k}<0$).
This paper is organized as follows: Section II describes some basic properties of FRBs. Section III presents an overview of distance calculations in cosmology within the FLRW framework. In Section IV we discuss the methodology used to reconstruct $H(z)$ and the comoving distance from FRB data, which are subsequently employed to infer the cosmic curvature density parameter. The data utilized in this work are detailed in Section V. Our results and discussion are presented in Section VI. Finally, Section VII contains our conclusions.
II Basic properties of Fast Radio Bursts
The observed dispersion measure, $\mathrm{DM}_{\mathrm{obs}}$, is the sum of local and extragalactic contributions, expressed as:
$\mathrm{DM}_{\mathrm{obs}}=\mathrm{DM}_{\mathrm{local}}+\mathrm{DM}_{\mathrm{EG}}(z),$ (1)
where the local component includes contributions from the Milky Way:
$\mathrm{DM}_{\mathrm{local}}=\mathrm{DM}_{\mathrm{ISM}}+\mathrm{DM}_{\mathrm{halo}},$ (2)
and the extragalactic component includes contributions from the intergalactic medium (IGM) and the host galaxy of FRBs:
$\mathrm{DM}_{\mathrm{EG}}=\mathrm{DM}_{\mathrm{IGM}}+\frac{\mathrm{DM}_{\mathrm{host}}}{(1+z)}.$ (3)
Each term in this decomposition corresponds to distinct astrophysical environments. $\mathrm{DM}_{\mathrm{ISM}}$ accounts for free electrons in the interstellar medium of the Milky Way and is commonly estimated using galactic electron density models like NE2001 Cordes and Lazio (2002) and YMW16 Yao et al. (2017). We adopt NE2001, as recent studies suggest that YMW16 may overestimate $\mathrm{DM}_{\mathrm{ISM}}$ at low Galactic latitudes Ocker et al. (2021). The Galactic halo contribution, $\mathrm{DM}_{\mathrm{halo}}$, is less constrained but estimated in the range $50-100~{}\mathrm{pc}\,\mathrm{cm}^{-3}$ Prochaska and Zheng (2019), based on observational constraints, such as the dispersion measure to the Large Magellanic Cloud (LMC), high-velocity cloud dynamics, and hydrostatic equilibrium models of the galactic halo gas. We then adopt a conservative value of $\mathrm{DM}_{\mathrm{halo}}=50~{}\mathrm{pc}\,\mathrm{cm}^{-3}$.
The intergalactic medium (IGM) dispersion measure, $\mathrm{DM}_{\mathrm{IGM}}$, dominates the observed $\mathrm{DM}_{\mathrm{obs}}$ and is crucial for cosmology. However, $\mathrm{DM}_{\mathrm{IGM}}$ exhibits significant scatter due to electron distribution inhomogeneities along the line of sight. Its mean value is given by Deng and Zhang (2014):
[…]
where $c$ is the speed of light, $G$ the gravitational constant, and $m_{p}$ the proton mass. $\Omega_{b}$ is the cosmic baryon density, and $E(z)=H(z)/H_{0}$ the normalized Hubble parameter. The term $f_{\mathrm{IGM}}(z)$ denotes the baryon fraction in the intergalactic medium, which we take as $f_{\mathrm{IGM}}=0.82\pm 0.04$ Shull et al. (2012); Zhou et al. (2014). The factor $f_{e}(z)$ describes the baryons ionization state and is defined as:
$f_{e}(z)=Y_{H}X_{e,H}(z)+\frac{1}{2}Y_{\mathrm{He}}X_{e,\mathrm{He}}(z),$ (5)
where $Y_{H}=0.75$ and $Y_{\mathrm{He}}=0.25$ are the mass fractions of hydrogen and helium, respectively. The terms $X_{e,H}(z)$ and $X_{e,\mathrm{He}}(z)$ denote their ionization fractions. Since both are fully ionized at $z<3$, we set $X_{e,H}=X_{e,\mathrm{He}}=1$.
The host galaxy contribution, $\mathrm{DM}_{\mathrm{host}}$, depends on the local environment of the FRB and varies with redshift due to cosmological expansion. Although some studies assume a constant value (for example, $\mathrm{DM}_{\mathrm{host}}=100~{}\mathrm{pc}\,\mathrm{cm}^{-3}$ Tendulkar et al. (2017)), this can yield unphysical results at low redshifts, including $\mathrm{DM}_{\mathrm{IGM}}$ negative values. To address this issue, we adopt the redshift-dependent parametrization proposed by Zhang et al. (2020), based on the IllustrisTNG simulation:
$\mathrm{DM}_{\mathrm{host}}=A(1+z)^{\alpha},$ (6)
where $A$ and $\alpha$ are free parameters that depend on the FRB type and host galaxy properties. For instance, repeating FRBs in spiral galaxies follow $\mathrm{DM}_{\mathrm{host}}=34.72(1+z)^{1.08}~{}\mathrm{pc}\,\mathrm{cm}^{-3}$, while non-repeating ones follow $\mathrm{DM}_{\mathrm{host}}=32.97(1+z)^{0.84}~{}\mathrm{pc}\,\mathrm{cm}^{-3}$.
The decomposition in Eq. (1) isolates $\mathrm{DM}_{\mathrm{IGM}}$, which is crucial for using FRBs as cosmological probes, as discussed in the following sections.
III Distances in cosmology
The comoving distance $D_{C}$ is a key cosmological quantity, providing an invariant measure of spatial separation between an observer and a distant object, independent of cosmological expansion. In the FLRW framework, $D_{C}$ is defined as:
$D_{C}(z)=c\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}.$ (7)
A related quantity is the horizon distance $D_{H}$, which represents the maximum scale over which information could have been exchanged since the Big Bang. It is defined as:
$D_{H}=\frac{c}{H_{0}}.$ (8)
This distance sets a fundamental scale in cosmology and is often used to normalize other cosmological distances.
On the other hand, the metric distance $D_{M}$ is related to the comoving distance and depends on the curvature. For different values of the curvature parameter $\Omega_{k}$, its functional form is Weinberg (1972):
$D_{M}=\begin{cases}\frac{D_{H}}{\sqrt{\Omega_{k}}}\sinh\left(\sqrt{\Omega_{k}}\frac{D_{C}}{D_{H}}\right),&\text{if }\Omega_{k}>0,\\ D_{C},&\text{if }\Omega_{k}=0,\\ \frac{D_{H}}{\sqrt{|\Omega_{k}|}}\sin\left(\sqrt{|\Omega_{k}|}\frac{D_{C}}{D_{H}}\right),&\text{if }\Omega_{k}<0,\end{cases}$ (9)
where $\Omega_{k}=0$ for a flat universe, $\Omega_{k}>0$ for an open universe and $\Omega_{k}<0$ for a closed universe.
In observational cosmology, absolute distance measurements are difficult to obtain; instead, dimensionless ratios normalized by the sound horizon $r_{d}$ are commonly used. Two ratios frequently used in BAO analysis are:
$\frac{D_{M}(z)}{r_{d}},\quad\text{and}\quad\frac{D_{H}(z)}{r_{d}}.$ (10)
The sound horizon $r_{d}$, is the maximum distance that acoustic waves could travel in the primordial plasma before photon decoupling and is given by Padilla et al. (2021):
$r_{d}=\int_{z_{d}}^{\infty}\frac{c_{s}dz}{H(z)},$ (11)
where $c_{s}$ is the speed of sound in the primordial plasma and $z_{d}$ denotes the baryon drag redshift, corresponding to the epoch when baryons decoupled from radiation. Measuring the scale $r_{d}$ from BAO data provides an independent way to test cosmological models and infer the universe’s expansion rate.
The quantities $D_{M}/r_{d}$ and $D_{H}/r_{d}$ serve as angular and radial standard rulers, respectively. $D_{M}/r_{d}$ determines the angular scale of BAO features in galaxy distributions, while $D_{H}/r_{d}$ determines the radial scale of the BAO peak in redshift-space distortions. BAO surveys measure the peak in both directions: perpendicular and along the line of sight, tracing $D_{M}/r_{d}$ and $D_{H}/r_{d}$, respectively. This dual measurement provides constraints on the expansion history $H(z)$ and the curvature parameter $\Omega_{k}$, offering an independent probe of the cosmological model.
In practice, the metric distance is not directly observable. Instead, we measure the angular diameter distance $D_{A}$. Due to cosmic expansion, $D_{A}$ is related to the comoving transverse distance $D_{M}$ by:
$D_{A}(z)=\frac{D_{M}(z)}{1+z}.$ (12)
By combining distance ratios $D_{M}/r_{d}$ and $r_{d}/D_{H}$, we obtain an expression for $D_{A}$ that depends only on observational quantities Gong et al. (2024); Liu et al. (2025):
$D_{A}(z)=\frac{c}{(1+z)\,H(z)}\,\frac{D_{M}}{r_{d}}\,\frac{r_{d}}{D_{H}}.$ (13)
This formulation is useful for inferring $D_{A}$ in a model-independent way, relying solely on measurements of the Hubble parameter and BAO-derived distance ratios. It enables cosmological tests without assuming a specific value of $r_{d}$, making it robust to uncertainties in early-universe physics. We will use this equation later to combine FRBs and BAO da in section VI.
IV Methodology
Following the discussions in the previous sections, we now describe the reconstruction of the comoving distance using real FRB data and its integration with BAO measurements. This section also outlines the technical procedures to derive angular diameter distances and constrain the curvature.
IV.1 Extracting $D_{C}$ and $D_{A}$ from $\langle\mathrm{DM_{IGM}}\rangle$
To infer the comoving distance from the dispersion measure, we rewrite the Macquart relation, i.e., Eq.(4), as:
[…]
with
$A=\frac{3c\,\Omega_{b}H_{0}^{2}}{8\pi Gm_{p}},$ (15)
where the free electron distribution in the IGM introduces an additional modulation described by the function:
$F(z)=(1+z)f_{\mathrm{IGM}}(z)f_{e}(z),$ (16)
and $\dot{D}_{C}(z)=dD_{C}(z)/dz$.
The redshift-dependent factor $F(z)$ must be removed from $\langle\mathrm{DM}_{\mathrm{IGM}}\rangle$. This is done by differentiating Eq. (14) with respect to redshift (note that Eq. (7) was also derived.)
$\dot{D_{C}}(z)=\frac{c}{H(z)}=\frac{1}{AF(z)}\frac{d\langle\mathrm{DM_{IGM}}(z)\rangle}{dz}\,,$ (17)
and then integrating $c/H(z)$, thus yielding a direct estimate of $D_{C}$ from the FRB data.
The key step in reconstructing $D_{C}$ is obtaining the derivative $d\langle\mathrm{DM_{IGM}(z)\rangle}\left(dz\right)^{-1}$. Although this method was first proposed in Yu and Wang (2017), the lack of a sufficient number of localized FRBs at the time led to the use of simulations and Gaussian Processes. In contrast, here we apply the method to real data, using artificial neural networks (ANNs) to reconstruct the derivative and constrain cosmic curvature. In particular, we use the method developed in Fortunato et al. (2025), in which ANNs are trained to reconstruct the relation between $\mathrm{DM_{IGM}}$ and $z$ in a data-driven manner.
Specifically, we adopted a Multilayer Perceptron (MLP) architecture implemented using the scikit-learn library Pedregosa et al. (2011), applying hyperparameter tuning via GridSearchCV to identify the optimal model. Various configurations of hidden layer sizes, activation functions, solvers, and learning rates were tested. The optimal model was selected based on cross-validation performance and learning curve analysis, ensuring that overfitting was minimized and that the network generalizes well to unseen data (see Fortunato et al. (2025) for more details). Bootstrap resampling was used to estimate uncertainties in the ANNs predictions, enabling robust error propagation to the reconstructed derivative $\frac{d\langle\mathrm{DM_{IGM}}\rangle}{dz}$ and, consequently, to $D_{C}(z)$.
Once $D_{C}(z)$ is known, the angular diameter distance can be derived solely from the FRB data using Eqs. (9) and (12). From now on, we refer to the $D_{A}(z)$ obtained exclusively from FRB data as $D^{\mathrm{FRB}}_{A}(z)$. It is important to emphasize that the determination of both $D_{C}(z)$ and $D^{\mathrm{FRB}}_{A}(z)$ is entirely independent of any cosmological model and is derived directly from the data.
IV.2 Estimation of the cosmic curvature
The key aspect of our approach to constraining spatial curvature is that FRBs play a dual role in the process. First, as described above, they allow us to directly obtain $D^{\mathrm{FRB}}_{A}(z)$ from the data. Second, as demonstrated in Fortunato et al. (2025), the quantity $\left(\frac{d\langle\mathrm{DM_{IGM}}(z)\rangle}{dz}\right)^{-1}$ enables the reconstruction of $H(z)$ and $H_{0}$ — see Eq. (17). The Hubble parameter derived from FRBs can also be combined with BAO measurements of the ratios $D_{M}/r_{d}$ and $D_{H}/r_{d}$, enabling the determination of the angular diameter distance via Eq. (13) in a cosmology-independent manner, using only the data. We shall refer to the angular diameter distance obtained in this way as $D^{\mathrm{BAO+FRB}}_{A}(z)$.
This approach allows us to compare $D^{\mathrm{FRB}}_{A}(z)$ and $D^{\mathrm{BAO+FRB}}_{A}(z)$ to estimate the spatial curvature parameter $\Omega_{k}$ in a data-driven manner. In both cases of $D_{A}$ estimation, the Hubble constant $H_{0}$ is not externally imposed but is self-consistently derived by extrapolating the FRB-based reconstruction of $H(z)$ to redshift $z=0$. This internal determination of $H_{0}$ helps mitigate the degeneracy between $H_{0}$ and $\Omega_{k}$, avoiding the need for external priors that could bias their coupling. Moreover, since $H_{0}$ enters both $D_{A}$ and $D_{C}$ symmetrically, variations in its value do not affect the determination of $\Omega_{k}$ within this framework.
To constrain $\Omega_{k}$, we implement two distinct approaches:
  • Covariance matrix approach: In our first approach, we account for the statistical correlation among the distance data points, primarily because both the angular diameter distance $D_{A}(z)$ and the comoving distance $D_{C}(z)$ depend on the same reconstructed Hubble parameter $H(z)$. To properly account for these correlations in the likelihood analysis, we construct a covariance matrix that incorporates both statistical uncertainties and correlated terms, following the strategy proposed in Liu et al. (2025). The likelihood function is given by
    $\ln\mathcal{L}=-\frac{1}{2}\Delta D_{i}^{T}\,\mathbf{Cov}^{-1}_{ij}\,\Delta D_{j},$ (18)
    where the residuals are defined as $\Delta D=D_{A}^{\mathrm{BAO+FRB}}(z_{i})-D_{A}^{\mathrm{FRB}}(z_{i};\Omega_{k})$, and the total covariance matrix is expressed as
    $\mathbf{Cov}_{ij}=\mathbf{Cov}^{\mathrm{stat}}_{ii}+\mathbf{Cov}^{\mathrm{corr}}_{ij}.$ (19)
    The statistical part captures the propagated uncertainties from the reconstruction of $H(z)$ and the BAO-based measurements. The correlated term reflects the intrinsic coupling between $D_{A}$ and $D_{C}$, since both quantities are derived from the same FRB-based estimates of $H(z)$. This term is computed using the standard definition of covariance: $\mathrm{Cov}(X_{i},Y_{j})=\mathbb{E}[(X_{i}-\mu_{i})(Y_{j}-\mu_{j})]$.
    Although the DESI collaboration provides a fiducial covariance matrix, it is derived under specific model assumptions and tailored to their reconstruction method. In contrast, our analysis relies on a model-independent reconstruction of $H(z)$ and therefore requires a covariance matrix that accurately reflects our methodology. By constructing the covariance matrix directly from our reconstructed distances, we ensure internal consistency and avoid potential mismatches. Moreover, the authors in Liu et al. (2025) have shown that this method yields results consistent with the DESI covariance estimates, further supporting the reliability of this approach.
  • Gaussian $\chi^{2}$ approach: For comparison, we also implement a simpler approach in which all data points are treated as independent and Gaussian-distributed. In this case, we use the traditional $\chi^{2}$ minimization method:
    $\chi^{2}=\sum_{i}\frac{\left[D_{A}^{\mathrm{BAO+FRB}}(z_{i})-D_{A}^{\mathrm{FRB}}(z_{i};\Omega_{k})\right]^{2}}{\sigma_{i}^{2}},$ (20)
    where $\sigma_{i}$ represents the combined error from the reconstruction of $D_{A}$ and $D_{C}$. This method does not account for correlations among the data points and may underestimate the uncertainty, but it serves as a useful baseline.
V Data
To implement our analysis, we use a subset of 80 from the 92 localized FRBs compiled by Wang et al. (2025). In addition to requiring known redshifts and dispersion measures, we impose a selection criterion also introduced in Wang et al. (2025) to ensure the physical consistency of the intergalactic medium contribution to the dispersion measure, $\mathrm{DM}_{\mathrm{IGM}}$. Specifically, we require
$\mathrm{DM}_{\mathrm{obs}}-\mathrm{DM}_{\mathrm{MW}}>80~{}\mathrm{pc~{}cm^{-3}},$
which accounts for a typical range of halo contributions, $\mathrm{DM}_{\mathrm{halo}}\sim 50\text{--}80~{}\mathrm{pc~{}cm^{-3}}$. This criterion serves as a practical proxy to enforce $\mathrm{DM}_{\mathrm{obs}}-\mathrm{DM}_{\mathrm{MW}}-\mathrm{DM}_{\mathrm{halo}}>0$, ensuring that the inferred $\mathrm{DM}_{\mathrm{IGM}}$ remains physically meaningful and avoids negative values. This cut also helps to eliminate FRBs with low total extragalactic dispersion, which could lead to biased inferences of the Hubble parameter, particularly by favouring artificially high values of $H_{0}$ within a narrow prior range, as shown in Wang et al. (2025). Furthermore, this selection indirectly filters out several low-redshift FRBs that may be heavily influenced by peculiar velocities of their host galaxies and the local cosmic web. By excluding these events, we minimize the impact of non-cosmological effects on the reconstruction of $H(z)$ and improve the robustness of the derived distances.
Data source $z$ $D_{M}/r_{d}$ $D_{H}/r_{d}$ Ref.
BOSS galaxy–galaxy 0.38 $10.23\pm 0.17$ $25.00\pm 0.76$ Alam et al. (2017)
eBOSS galaxy–galaxy 0.51 $13.36\pm 0.21$ $22.33\pm 0.58$ Alam et al. (2017)
eBOSS galaxy–galaxy 0.70 $17.86\pm 0.33$ $19.33\pm 0.62$ Bautista et al. (2021); Gil-Marín et al. (2020)
eBOSS galaxy–galaxy 0.85 $19.60\pm 0.41$ $19.60\pm 0.72$ Tamone et al. (2020); De Mattia et al. (2021)
eBOSS galaxy–galaxy 1.48 $30.21\pm 0.79$ $13.23\pm 0.47$ Neveux et al. (2020); Hou et al. (2021)
DESI DR2 LRG1 0.510 $13.588\pm 0.167$ $21.863\pm 0.425$ Karim et al. (2025)
DESI DR2 LRG2 0.706 $17.351\pm 0.177$ $19.455\pm 0.330$ Karim et al. (2025)
DESI DR2 LRG3+ELG1 0.934 $21.576\pm 0.152$ $17.641\pm 0.193$ Karim et al. (2025)
DESI DR2 ELG2 1.321 $27.601\pm 0.318$ $14.176\pm 0.221$ Karim et al. (2025)
DESI DR2 QSO 1.484 $30.512\pm 0.760$ $12.817\pm 0.516$ Karim et al. (2025)
Table 1: Measurements of $D_{M}/r_{d}$ and $D_{H}/r_{d}$ from various surveys.
In addition, seven events were excluded from our analysis for specific reasons. First, FRB200110E is omitted because it is located at an extremely close distance of only $3.6~{}\mathrm{Mpc}$, within a globular cluster in the M81 galaxy Kirsten et al. (2022); Bhardwaj et al. (2021). Its proximity implies a negligible contribution from the intergalactic medium, making it unsuitable for cosmological inference. Second, FRB210117 and FRB190520 display dispersion measures significantly exceeding theoretical expectations based on their redshift. These anomalies classify them as statistical outliers, probably influenced by a dense or highly magnetized local environment in their host galaxy Bhandari et al. (2023); Gordon et al. (2023); Niu et al. (2022). Moreover, we did not reconstruct the functions at higher redshift values due to the lack of FRB data points in this regime. Although there are four points around $z\sim 1$ — FRB220610, FRB221029, FRB230521, FRB240123 — we excluded them because they could bias the reconstruction, given the gap between them and the last point used in this study at $z=0.66$. The remaining 80 FRBs form a clean and physically consistent dataset for the curvature analysis presented in the following sections.
For the analysis that follows, we employ a total of ten BAO data points in the range of the reconstructed FRBs redshifts, z = [0, 1.5]: five from the BOSS/eBOSS programs and five from the recent DESI DR2 release, summarized in Table 1. These measurements are used in combination with our FRB and ANN-based reconstructions of $H(z)$ and distances, to constrain the cosmic curvature parameter $\Omega_{k}$.
VI Results and Discussions
To apply the methodology described above, we first subtract the contributions from the Milky Way, the halo, and the host galaxy in order to obtain $\mathrm{DM_{IGM}}$ from the observed dispersion measures of the subset of 80 FRBs described in Sec. V. Then we reconstruct the comoving distance $D_{C}(z)$. The first step involves determining the redshift evolution of the average intergalactic dispersion measure, $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$. As detailed in Sec. IV, this is performed using an ANN algorithm (see Fortunato et al. (2025) for more details). The result is shown in panel (a) in Figure 1. The red points with error bars represent the individual FRB measurements and their associated uncertainties, while the solid black curve corresponds to the ANN-based reconstruction. The shaded region around it denotes the $1\sigma$ confidence interval.
Figure 1: Reconstructed FRBs observables. Left: redshift evolution of the average IGM dispersion measure $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$. Right: comoving distance $D_{C}(z)$ inferred from FRBs.
Once $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$ is reconstructed, it is used to derive the comoving distance $D_{C}(z)$ (see Eq. (17)). The resulting comoving distance curve is displayed in panel (b) in Figure 1. The dashed black line shows the reconstructed $D_{C}(z)$, while the shaded blue band represents the associated $1\sigma$ uncertainty. This reconstruction provides a continuous cosmological model-independent estimate of the comoving distance across a range of redshifts, from low redshifts up to $z\sim 1.5$, where BAO measurements are also available for comparison. Subsequently, $D^{\mathrm{FRB}}_{A}(z)$ and $D^{\mathrm{BAO+FRB}}_{A}(z)$ are computed.
To constrain $\Omega_{k}$, we explore the likelihood using the emcee Python module, a Markov Chain Monte Carlo (MCMC) ensemble sampler Foreman-Mackey et al. (2013). The prior for the Hubble constant is taken as the extrapolated value $H_{0}=H(z=0)$, obtained from the ANN reconstruction of the FRB data. The final results from both approaches (covariance matrix and $\chi^{2}$ minimization) are presented and compared in Table 2, and the corresponding posterior distributions are shown in Figure 2. These results demonstrate that the combined use of FRBs and BAO data provides a robust and independent constraint on the spatial geometry of the Universe.
Figure 2: Posterior probability distributions for the cosmic curvature parameter $\Omega_{k}$ under two assumptions. Left: using the full covariance matrix Right: Neglecting correlations.
Survey $\Omega_{k}$ (with covariance) $\Omega_{k}$ (without covariance)
eBOSS $-0.25^{+0.35}_{-0.46}$ $-0.18\pm 0.20$
DESI DR2 $0.10^{+0.33}_{-0.38}$ $-0.19\pm 0.08$
eBOSS + DESI DR2 $-0.09^{+0.24}_{-0.31}$ $-0.17\pm 0.06$
Table 2: Estimated values of the cosmic curvature parameter $\Omega_{k}$ using BAO and FRB comoving distance measurements.
From the covariance-based analysis (Table 2), we derive the following constraints:
  • The eBOSS data yield $\Omega_{k}=-0.25^{+0.35}_{-0.46}$, indicating a mild preference for a closed Universe, albeit with substantial uncertainties.
  • The DESI data alone provide a constraint centered around spatial flatness, with $\Omega_{k}=0.10^{+0.33}_{-0.38}$.
  • When combining the eBOSS and DESI DR2 datasets, the inferred curvature is $\Omega_{k}=-0.09^{+0.24}_{-0.31}$, suggesting a slightly closed Universe with reduced uncertainties compared to individual datasets.
In contrast, neglecting the off-diagonal elements of the covariance matrix (Table 2) yields significantly tighter posterior distributions:
  • For eBOSS, we find $\Omega_{k}=-0.18\pm 0.20$,
  • For DESI, $\Omega_{k}=-0.19\pm 0.08$, and
  • For the joint eBOSS + DESI DR2 analysis, $\Omega_{k}=-0.17\pm 0.06$.
Despite variations in central values, all estimates remain consistent with a spatially flat Universe ($\Omega_{k}=0$) at the $1\sigma$ level. The integration of FRB-based comoving distance reconstruction serves as an independent and complementary cosmological probe, reinforcing the growing utility of FRBs in precision cosmology.
Taken together, these results demonstrate that ANN reconstructions of comoving distances from FRBs, when combined with state-of-the-art BAO measurements, provide competitive constraints on spatial curvature, comparable to those obtained from other late-time cosmological probes Liu et al. (2025); Gong et al. (2024); Mukherjee and Banerjee (2022); Wang et al. (2021). While Gaussian-based methods yield tighter intervals, they may underestimate uncertainties due to neglected correlations. The apparent preference for negative values in this case may result from omitting the covariance matrix, which can bias the final result. In contrast, the approach that incorporates the full covariance matrix, though yielding broader confidence regions, provides a more conservative and robust assessment of parameter uncertainties. These findings highlight the critical role of accounting for the full covariance structure in cosmological parameter inference, especially when combining different datasets or when intrinsic correlations are present.
[…]
VII Conclusions
In this work, we presented a fully data-driven strategy to constrain the spatial curvature of the Universe by leveraging the synergy between FRBs and BAO measurements. Reconstructing the Hubble parameter $H(z)$ using an ANN trained on a set of localized FRBs, we derived the comoving distance $D_{C}(z)$ and subsequently estimated $D_{A}(z)$ to compare with the same quantity obtained from BAO+FRB data. This dual-path approach enabled robust constraints on the curvature parameter $\Omega_{k}$, independent of early-universe assumptions or fiducial cosmological models.
Our results indicate that, within current observational limits, all estimates remain consistent with a spatially flat Universe. However, a persistent mild preference for negative curvature is observed across multiple statistical treatments, suggesting that future improvements in data quantity and quality could be decisive. Incorporating the full covariance matrix in the likelihood analysis yields more conservative yet reliable uncertainty estimates, underscoring the importance of properly accounting for statistical correlations in precision cosmology.
Looking ahead, enhancing the constraining power of FRB-based curvature measurements will require advances on both observational and methodological fronts. Increasing the number of localized FRBs with secure redshift determinations will reduce statistical uncertainties and improve redshift coverage. The number of FRBs detected is indeed rapidly increasing with the advent of large-scale radio surveys and real-time localization capabilities, such as the Canadian Hydrogen Intensity Mapping Experiment (CHIME) Amiri et al. (2022), the Australian Square Kilometre Array Pathfinder (ASKAP) Qiu et al. (2023), the improved version of the Karoo Array Telescope (MeerKAT) Tian et al. (2024). At the same time, improved modeling of host galaxy and halo contributions to the dispersion measure will mitigate systematic errors in the extraction of $\mathrm{DM}_{\mathrm{IGM}}$. On the analytical side, incorporating hybrid inference techniques that combine neural networks with advanced statistical frameworks may enhance the reconstruction accuracy and stability.
As new FRBs discoveries expand the redshift baseline and improve the resolution of the available datasets, their integration with BAO measurements promises to yield increasingly precise and independent insights into the geometry of the cosmos. This work reinforces the case for FRBs as a valuable cosmological probe and sets the stage for their broader inclusion in future large-scale structure analyses.
Acknowledgements.
JASF thanks FAPES for their financial support. WSHR acknowledges partial financial support from FAPES. G.E.R. acknowledges financial support from the State Agency for Research of the Spanish Ministry of Science and Innovation under grant PID2022-136828NB-C41/AEI/10.13039/501100011033/, and by “ERDF A way of making Europe”, by the “European Union”. He also thanks support from PIP 0554 (CONICET).

### v2

@@
 The spatial curvature of the Universe remains a central question in modern cosmology.
-In this work, we explore the potential of localized Fast Radio Bursts (FRBs) as a novel tool to constrain the cosmic curvature parameter $\Omega_{k}$ in a cosmological model-independent way.
-Using a sample of 80 FRBs with known redshifts and dispersion measures, we reconstruct the Hubble parameter $H(z)$ via artificial neural networks, which is then used to obtain angular diameter distances $D_{A}(z)$ through two complementary approaches.
-First, we derive the comoving distance $D_{C}(z)$ and $D_{A}(z)$ from FRBs without assuming a fiducial cosmological model.
-Then, the $H(z)$ reconstruction is combined with Baryon Acoustic Oscillation (BAO) measurements to infer $D_{A}(z)$.
-By comparing the FRB-derived and BAO+FRB-derived $D_{A}(z)$ values, we extract constraints on $\Omega_{k}$ under the Friedmann–Lemaître–Robertson–Walker metric.
-Our results show consistency with a spatially flat Universe within $1\sigma$ uncertainties, although a mild preference for negative curvature is observed across both covariance-based and Gaussian analyses.
-This study highlights the growing relevance of FRBs in precision cosmology and demonstrates their synergy with BAO data as a powerful probe of the large-scale geometry of the Universe.
+In this work, we explore the potential of localized Fast Radio Bursts (FRBs) as a novel tool to constrain the cosmic curvature parameter $\Omega_{k}$ in a cosmological model–independent way.
+Using a sample of 120 FRBs with known redshifts and dispersion measures, we reconstruct the Hubble parameter $H(z)$ via artificial neural networks, and use it to obtain angular–diameter distances $D_{A}(z)$ through two complementary approaches.
+First, we derive the comoving distance $D_{C}(z)$ and $D_{A}(z)$ directly from FRBs without assuming a fiducial cosmology.
+Then, we combine the FRB-based $H(z)$ with Baryon Acoustic Oscillation (BAO) DESI DR2 measurements to infer $D_{A}(z)$.
+By comparing the FRB-derived and BAO+FRB-derived $D_{A}(z)$, we constrain spatial curvature.
+Our covariance-based likelihood (accounting for correlated uncertainties) yields $\Omega_{k}=-0.31\pm 0.57$, while a diagonal (Gaussian) treatment gives $\Omega_{k}=-0.13\pm 0.46$.
+Both estimations are consistent with spatial flatness at the $1\sigma$ level, albeit with a mild preference for negative curvature.
+Explicitly accounting for the full covariance broadens the intervals and avoids underestimation of uncertainties.
+These results highlight the growing relevance of FRBs in precision cosmology and their synergy with BAO as a powerful, cosmological model–independent probe of the large–scale geometry of the Universe.
 Keywords: Fast Radio Bursts, Artificial Neural Networks, Cosmic Distances, Curvature of Universe.
@@
 To obtain this estimate, they used a combination of datasets, including BAO, Cosmic Chronometers (CC), and Strong Gravitational Lensing (SGL).
-However, it is important to recognize that these constraints are inherently cosmological model-dependent, relying on assumptions within the non-flat $\Lambda$CDM framework, which suffers from the well-known intrinsic degeneracy between cosmic curvature and the Hubble constant.
-Moreover, as with the Hubble tension, the apparent disagreement in curvature estimates raises the question of whether systematic biases or new physics are at play.
-In this context, exploring methods that provide a complementary and cosmological model-independent perspective is particularly valuable.
-Developing curvature measurements that are independent of the cosmological model, especially those that do not rely solely on CMB-based assumptions, offers a promising pathway to gaining a better understanding of these issues.
+Other recent studies reached conclusions consistent with spatial flatness by combining alternative datasets, reporting mean values of $\Omega_{k}$ that deviate slightly from zero but remain within $1\sigma$ of flatness.
+For instance, in Vagnozzi et al.
+(2021a) used full-shape galaxy power spectra from BOSS DR12 CMASS and Planck data, finding $\Omega_{k}=0.0023\pm 0.0028$, while Vagnozzi et al.
+(2021b) combined Planck with CC data to obtain $\Omega_{k}=-0.0054\pm 0.0055$.
+[…]
 Recent studies Gong et al.
@@
 (2022), constrain dark energy equation of state Zhou et al.
-(2014); Gao et al.
+(2014a); Gao et al.
 (2014), trace the magnetic fields in the IGM Akahori et al.
@@
 Consequently, the background spacetime is described by the Friedmann–Lemaître–Robertson–Walker (FLRW) metric, with the curvature parameter $\Omega_{k}$ determining the geometry of the Universe: flat ($\Omega_{k}=0$), open ($\Omega_{k}>0$), or closed ($\Omega_{k}<0$).
+Throughout this work, the expression model-independent denotes that no early-universe calibration or specific cosmological model is assumed: BAO enter only through dimensionless ratios whose combination cancels $r_{d}$, and $H(z)$ is reconstructed directly from FRB data via the Macquart relation.
+This approach relies on Eqs.
+(4) and (5) (defined in Sec.
+II), which connect the FRB dispersion measures to $H(z)$ via minimal physical assumptions (FLRW geometry and standard ionization fractions) without assuming a particular cosmological model.
+Curvature effects enter explicitly only in the FRB-only through the $S_{k}$ function mapping from $D_{C}$ to $D_{M}$.
 This paper is organized as follows: Section II describes some basic properties of FRBs.
@@
 Its mean value is given by Deng and Zhang (2014):
-[…]
+$\langle\mathrm{DM}_{\mathrm{IGM}}\rangle=\left(\frac{3c}{8\pi Gm_{p}}\right)\Omega_{b}H_{0}\int_{0}^{z}\frac{(1+z^{\prime})f_{\mathrm{IGM}}(z^{\prime})f_{e}(z^{\prime})}{E(z^{\prime})}dz^{\prime},$ (4)
 where $c$ is the speed of light, $G$ the gravitational constant, and $m_{p}$ the proton mass.
@@
 (2012); Zhou et al.
-(2014).
+(2014b).
 The factor $f_{e}(z)$ describes the baryons ionization state and is defined as:
@@
 $D_{C}(z)=c\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}.$ (7)
-A related quantity is the horizon distance $D_{H}$, which represents the maximum scale over which information could have been exchanged since the Big Bang. It is defined as:
-$D_{H}=\frac{c}{H_{0}}.$ (8)
-This distance sets a fundamental scale in cosmology and is often used to normalize other cosmological distances.
+A related scale is the Hubble distance (or Hubble radius) at the present epoch,
+$D_{H_{0}}=\frac{c}{H_{0}},$ (8)
+with the general redshift-dependent definition
+$D_{H}(z)=\frac{c}{H(z)}.$ (9)
+This scale is often used to normalize cosmological distances and should not be confused with causal horizons (particle or event horizons), which involve time integrals of $H(z)$.
 On the other hand, the metric distance $D_{M}$ is related to the comoving distance and depends on the curvature.
 For different values of the curvature parameter $\Omega_{k}$, its functional form is Weinberg (1972):
-$D_{M}=\begin{cases}\frac{D_{H}}{\sqrt{\Omega_{k}}}\sinh\left(\sqrt{\Omega_{k}}\frac{D_{C}}{D_{H}}\right),&\text{if }\Omega_{k}>0,\\ D_{C},&\text{if }\Omega_{k}=0,\\ \frac{D_{H}}{\sqrt{|\Omega_{k}|}}\sin\left(\sqrt{|\Omega_{k}|}\frac{D_{C}}{D_{H}}\right),&\text{if }\Omega_{k}<0,\end{cases}$ (9)
+$D_{M}=\begin{cases}\frac{D_{H_{0}}}{\sqrt{\Omega_{k}}}\sinh\left(\sqrt{\Omega_{k}}\frac{D_{C}}{D_{H_{0}}}\right),&\text{if }\Omega_{k}>0,\\ D_{C},&\text{if }\Omega_{k}=0,\\ \frac{D_{H_{0}}}{\sqrt{|\Omega_{k}|}}\sin\left(\sqrt{|\Omega_{k}|}\frac{D_{C}}{D_{H_{0}}}\right),&\text{if }\Omega_{k}<0,\end{cases}$ (10)
 where $\Omega_{k}=0$ for a flat universe, $\Omega_{k}>0$ for an open universe and $\Omega_{k}<0$ for a closed universe.
@@
 $D_{A}(z)=\frac{c}{(1+z)\,H(z)}\,\frac{D_{M}}{r_{d}}\,\frac{r_{d}}{D_{H}}.$ (14)
+The equation above provides an angular diameter distance estimate that combines BAO ratios with an FRB-based reconstruction of the Hubble parameter.
+In this construction, the curvature parameter $\Omega_{k}$ does not enter explicitly, since the BAO measurements appear only through the ratio of their reported quantities and the sound horizon cancels.
+By contrast, the FRB-only approach reconstructs the comoving distance from the dispersion measure and then maps it into a transverse comoving distance using Eq.
+(9), which explicitly involves $\Omega_{k}$.
+Thus, curvature dependence arises only in the FRB-only estimator, while the BAO+FRB estimator remains algebraically independent of $\Omega_{k}$.
 This formulation is useful for inferring $D_{A}$ in a model-independent way, relying solely on measurements of the Hubble parameter and BAO-derived distance ratios.
@@
 To infer the comoving distance from the dispersion measure, we rewrite the Macquart relation, i.e., Eq.(4), as:
-[…]
+$\langle\mathrm{DM_{IGM}}(z)\rangle=A\int_{0}^{z}F(z^{\prime})\,\dot{D}_{C}(z^{\prime})dz^{\prime},$ (15)
 with
@@
 Covariance matrix approach: In our first approach, we account for the statistical correlation among the distance data points, primarily because both the angular diameter distance $D_{A}(z)$ and the comoving distance $D_{C}(z)$ depend on the same reconstructed Hubble parameter $H(z)$.
-To properly account for these correlations in the likelihood analysis, we construct a covariance matrix that incorporates both statistical uncertainties and correlated terms, following the strategy proposed in Liu et al.
-(2025).
+To properly account for these correlations in the likelihood analysis, we construct a covariance matrix that incorporates both statistical uncertainties and correlated terms, following the strategy found in Liu et al.
+(2025), based on the error propagation theory Owen (2000).
 The likelihood function is given by
@@
 $\mathbf{Cov}_{ij}=\mathbf{Cov}^{\mathrm{stat}}_{ii}+\mathbf{Cov}^{\mathrm{corr}}_{ij}.$ (20)
-    The statistical part captures the propagated uncertainties from the reconstruction of $H(z)$ and the BAO-based measurements.
-The correlated term reflects the intrinsic coupling between $D_{A}$ and $D_{C}$, since both quantities are derived from the same FRB-based estimates of $H(z)$.
-This term is computed using the standard definition of covariance: $\mathrm{Cov}(X_{i},Y_{j})=\mathbb{E}[(X_{i}-\mu_{i})(Y_{j}-\mu_{j})]$.
+The DESI DR2 survey Karim et al.
+(2025) provides the covariance of the primary BAO observables $\mathbf{x}\equiv\big(D_{M}/r_{d},\;D_{H}/r_{d}\big)$ at each redshift, with a full (off–diagonal) covariance matrix $\mathbf{C}_{x}$.
+Since our analysis uses Eq.
+(14), we propagate BAO errors to $D_{A}$ via the Jacobian Owen (2000) $\mathbf{J}_{i}=\big(\partial f/\partial(D_{M}/r_{d}),\;\partial f/\partial(D_{H}/r_{d})\big)\big|_{z_{i}}$:
+$\mathbf{C}_{D_{A}}^{\mathrm{(BAO)}}=\mathbf{J}\,\mathbf{C}_{x}\,\mathbf{J}^{\!\top}.$ (21)
+Following Eq. (20), only its diagonal enters $\mathbf{Cov}^{\mathrm{stat}}$, and we add in quadrature the FRB-inferred $H(z)$ contribution on the data side:
+$\big(\mathbf{Cov}^{\mathrm{stat}}\big)_{ii}=\big(\mathbf{C}_{D_{A}}^{\mathrm{(BAO)}}\big)_{ii}+\left(\frac{\partial D_{A}}{\partial H}\sigma_{H}\right)^{2}_{z_{i}}.$ (22)
+The correlated term captures the intrinsic coupling between $D_{A}^{\mathrm{FRB}}$ and $D_{C}^{\mathrm{FRB}}$ induced by the same FRB–reconstructed $H(z)$.
+We estimate it empirically from the FRB ensembles (bootstrap realizations) as a cross–covariance:
+$\mathbf{Cov}^{\mathrm{corr}}_{ij}$ $=$ $\mathrm{Cov}\!\left[D_{A}^{\mathrm{FRB}}(z_{i}),\,D_{C}^{\mathrm{FRB}}(z_{j})\right]$ (23)
+$=$ $\mathbb{E}\!\left[\big(D_{A}^{\mathrm{FRB}}-\mu_{A}\big)_{i}\,\big(D_{C}^{\mathrm{FRB}}-\mu_{C}\big)_{j}\right].$ (24)
 Although the DESI collaboration provides a fiducial covariance matrix, it is derived under specific model assumptions and tailored to their reconstruction method.
@@
 By constructing the covariance matrix directly from our reconstructed distances, we ensure internal consistency and avoid potential mismatches.
-Moreover, the authors in Liu et al.
-(2025) have shown that this method yields results consistent with the DESI covariance estimates, further supporting the reliability of this approach.
+As a cross-check, we repeated the inference using the DESI-provided covariance matrix (projected in the same way) and obtained posteriors for $\Omega_{k}$ that were statistically indistinguishable from our baseline – central values and uncertainties changed negligibly.
 Gaussian $\chi^{2}$ approach: For comparison, we also implement a simpler approach in which all data points are treated as independent and Gaussian-distributed.
@@
 This method does not account for correlations among the data points and may underestimate the uncertainty, but it serves as a useful baseline.
+The likelihood therefore compares a curvature–independent quantity, defined in Eq.
+(13), with a curvature–dependent one obtained via Eq.
+(9), and the sensitivity to $\Omega_{k}$ emerges solely from this comparison.
+In summary, the COV analysis uses the full matrix $\mathbf{Cov}=\mathrm{diag}\!\big(\mathbf{Cov}^{\mathrm{stat}}\big)+\mathbf{Cov}^{\mathrm{corr}}$, where the BAO contribution to $D_{A}$ is rigorously propagated from DESI’s (off–diagonal) covariance via the Jacobian, while the correlated FRB term is measured from the FRB ensemble.
+For comparison, the Gaussian case adopts only the diagonal variance $\sigma_{i}^{2}=\big(\mathbf{C}_{D_{A}}^{\mathrm{(BAO)}}\big)_{ii}+\left(\partial D_{A}/\partial H\,\sigma_{H}\right)^{2}_{z_{i}}$, i.e.
+it neglects $\mathbf{Cov}^{\mathrm{corr}}$.
 V Data
-To implement our analysis, we use a subset of 80 from the 92 localized FRBs compiled by Wang et al.
-(2025).
+To implement our analysis, we use a subset of 120 from the 131 localized FRBs presented in table LABEL:tab:frb_sample.
 In addition to requiring known redshifts and dispersion measures, we impose a selection criterion also introduced in Wang et al.
@@
 By excluding these events, we minimize the impact of non-cosmological effects on the reconstruction of $H(z)$ and improve the robustness of the derived distances.
-Data source $z$ $D_{M}/r_{d}$ $D_{H}/r_{d}$ Ref.
-BOSS galaxy–galaxy 0.38 $10.23\pm 0.17$ $25.00\pm 0.76$ Alam et al. (2017)
-eBOSS galaxy–galaxy 0.51 $13.36\pm 0.21$ $22.33\pm 0.58$ Alam et al. (2017)
-eBOSS galaxy–galaxy 0.70 $17.86\pm 0.33$ $19.33\pm 0.62$ Bautista et al. (2021); Gil-Marín et al. (2020)
-eBOSS galaxy–galaxy 0.85 $19.60\pm 0.41$ $19.60\pm 0.72$ Tamone et al. (2020); De Mattia et al. (2021)
-eBOSS galaxy–galaxy 1.48 $30.21\pm 0.79$ $13.23\pm 0.47$ Neveux et al. (2020); Hou et al. (2021)
-DESI DR2 LRG1 0.510 $13.588\pm 0.167$ $21.863\pm 0.425$ Karim et al. (2025)
-DESI DR2 LRG2 0.706 $17.351\pm 0.177$ $19.455\pm 0.330$ Karim et al. (2025)
-DESI DR2 LRG3+ELG1 0.934 $21.576\pm 0.152$ $17.641\pm 0.193$ Karim et al. (2025)
-DESI DR2 ELG2 1.321 $27.601\pm 0.318$ $14.176\pm 0.221$ Karim et al. (2025)
-DESI DR2 QSO 1.484 $30.512\pm 0.760$ $12.817\pm 0.516$ Karim et al. (2025)
-Table 1: Measurements of $D_{M}/r_{d}$ and $D_{H}/r_{d}$ from various surveys.
+Tracer $z$ $D_{M}/r_{d}$ $D_{H}/r_{d}$ Ref.
+LRG1 0.510 $13.588\pm 0.167$ $21.863\pm 0.425$ Karim et al. (2025)
+LRG2 0.706 $17.351\pm 0.177$ $19.455\pm 0.330$ Karim et al. (2025)
+LRG3+ELG1 0.934 $21.576\pm 0.152$ $17.641\pm 0.193$ Karim et al. (2025)
+ELG2 1.321 $27.601\pm 0.318$ $14.176\pm 0.221$ Karim et al. (2025)
+QSO 1.484 $30.512\pm 0.760$ $12.817\pm 0.516$ Karim et al. (2025)
+Lya 2.330 $338.988\pm 0.531$ $8.632\pm 0.101$ Karim et al. (2025)
+Table 1: Measurements of $D_{M}/r_{d}$ and $D_{H}/r_{d}$ from DESI DR2 survey.
 In addition, seven events were excluded from our analysis for specific reasons.
@@
 (2022).
-Moreover, we did not reconstruct the functions at higher redshift values due to the lack of FRB data points in this regime.
-Although there are four points around $z\sim 1$ — FRB220610, FRB221029, FRB230521, FRB240123 — we excluded them because they could bias the reconstruction, given the gap between them and the last point used in this study at $z=0.66$.
-The remaining 80 FRBs form a clean and physically consistent dataset for the curvature analysis presented in the following sections.
+A practical limitation of our reconstruction is the sparse sampling at high redshift ($z\gtrsim 1$).
+Our baseline analysis includes all available localized FRBs after cleaning the dataset (120 events in total), among which only a handful lie near or above $z\sim 1$ (notably FRB220610, FRB221029, FRB230521, FRB240123, and the most distant FRB240304B at $z=2.148$).
+In this regime, the ANN is weakly constrained by the data, which naturally broadens the predictive variance and, consequently, inflates the propagated uncertainties in $H(z)$, $D_{C}(z)$, and $D_{A}^{\mathrm{FRB}}(z)$.
+This effect is expected and reflects limited leverage rather than a modeling bias.
+Since our curvature inference is most strongly informed by the redshift range where BAO measurements exist ($z\lesssim 2.5$), the net impact is a conservative widening of error bands at high $z$ without qualitatively altering the central trends used to constrain $\Omega_{k}$.
 For the analysis that follows, we employ a total of ten BAO data points in the range of the reconstructed FRBs redshifts, z = [0, 1.5]: five from the BOSS/eBOSS programs and five from the recent DESI DR2 release, summarized in Table 1.
@@
 VI Results and Discussions
-To apply the methodology described above, we first subtract the contributions from the Milky Way, the halo, and the host galaxy in order to obtain $\mathrm{DM_{IGM}}$ from the observed dispersion measures of the subset of 80 FRBs described in Sec.
+To apply the methodology described above, we first subtract the contributions from the Milky Way, the halo, and the host galaxy in order to obtain $\mathrm{DM_{IGM}}$ from the observed dispersion measures of the subset of 120 FRBs described in Sec.
 V.
@@
 The red points with error bars represent the individual FRB measurements and their associated uncertainties, while the solid black curve corresponds to the ANN-based reconstruction.
-The shaded region around it denotes the $1\sigma$ confidence interval.
+The shaded region around it denotes the $3\sigma$ confidence interval.
 Figure 1: Reconstructed FRBs observables.
 Left: redshift evolution of the average IGM dispersion measure $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$.
-Right: comoving distance $D_{C}(z)$ inferred from FRBs.
-Once $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$ is reconstructed, it is used to derive the comoving distance $D_{C}(z)$ (see Eq.
+Right: Hubble parameter $H(z)$ inferred from FRBs.
+Once $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$ is reconstructed, it is used to derive the Hubble parameter $H(z)$ and by extrapolating the reconstruction to $z=0$ we obtain $H_{0}=68.74\pm 6.11\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$.
+Panel (b) of Figure 1 shows the reconstructed $H(z)$ curve with $3\sigma$ bands.
+This value is then applied throughout our methodology as described in Section IV to recover the comoving distance $D_{C}(z)$ (see Eq.
 (18)).
-The resulting comoving distance curve is displayed in panel (b) in Figure 1.
-The dashed black line shows the reconstructed $D_{C}(z)$, while the shaded blue band represents the associated $1\sigma$ uncertainty.
-This reconstruction provides a continuous cosmological model-independent estimate of the comoving distance across a range of redshifts, from low redshifts up to $z\sim 1.5$, where BAO measurements are also available for comparison.
+The resulting comoving distance curve is displayed in Figure 2.
+The dashed black line shows the reconstructed $D_{C}(z)$, while the shaded blue band represents the associated $3\sigma$ uncertainty.
+This reconstruction provides a continuous cosmological model-independent estimate of the comoving distance across a range of redshifts, from low redshifts up to $z\sim 2.5$, where BAO measurements are also available for comparison.
 Subsequently, $D^{\mathrm{FRB}}_{A}(z)$ and $D^{\mathrm{BAO+FRB}}_{A}(z)$ are computed.
+Figure 2: Comoving distance $D_{C}(z)$ reconstructed from FRBs.
 To constrain $\Omega_{k}$, we explore the likelihood using the emcee Python module, a Markov Chain Monte Carlo (MCMC) ensemble sampler Foreman-Mackey et al.
@@
 The prior for the Hubble constant is taken as the extrapolated value $H_{0}=H(z=0)$, obtained from the ANN reconstruction of the FRB data.
-The final results from both approaches (covariance matrix and $\chi^{2}$ minimization) are presented and compared in Table 2, and the corresponding posterior distributions are shown in Figure 2.
+The final results from both approaches (covariance matrix and Gaussian $\chi^{2}$ minimization) are presented in Figure 3, panel (a).
 These results demonstrate that the combined use of FRBs and BAO data provides a robust and independent constraint on the spatial geometry of the Universe.
-Figure 2: Posterior probability distributions for the cosmic curvature parameter $\Omega_{k}$ under two assumptions. Left: using the full covariance matrix Right: Neglecting correlations.
-Survey $\Omega_{k}$ (with covariance) $\Omega_{k}$ (without covariance)
-eBOSS $-0.25^{+0.35}_{-0.46}$ $-0.18\pm 0.20$
-DESI DR2 $0.10^{+0.33}_{-0.38}$ $-0.19\pm 0.08$
-eBOSS + DESI DR2 $-0.09^{+0.24}_{-0.31}$ $-0.17\pm 0.06$
-Table 2: Estimated values of the cosmic curvature parameter $\Omega_{k}$ using BAO and FRB comoving distance measurements.
-From the covariance-based analysis (Table 2), we derive the following constraints:
-  • The eBOSS data yield $\Omega_{k}=-0.25^{+0.35}_{-0.46}$, indicating a mild preference for a closed Universe, albeit with substantial uncertainties.
-  • The DESI data alone provide a constraint centered around spatial flatness, with $\Omega_{k}=0.10^{+0.33}_{-0.38}$.
-  • When combining the eBOSS and DESI DR2 datasets, the inferred curvature is $\Omega_{k}=-0.09^{+0.24}_{-0.31}$, suggesting a slightly closed Universe with reduced uncertainties compared to individual datasets.
-In contrast, neglecting the off-diagonal elements of the covariance matrix (Table 2) yields significantly tighter posterior distributions:
-  • For eBOSS, we find $\Omega_{k}=-0.18\pm 0.20$,
-  • For DESI, $\Omega_{k}=-0.19\pm 0.08$, and
-  • For the joint eBOSS + DESI DR2 analysis, $\Omega_{k}=-0.17\pm 0.06$.
+Figure 3: Left: One–dimensional posterior for the spatial curvature parameter $\Omega_{k}$ obtained with our two likelihoods.
+The red filled curve (COV) uses the full covariance between the FRB–derived $D_{C}$ and $D_{A}$ together with the BAO error propagation; the blue filled curve (Gaussian) adopts a diagonal treatment.
+The vertical dashed line marks $\Omega_{k}=0$.
+Right: Black points with $3\sigma$ bars are $D_{A}^{\rm BAO+FRB}(z)$ from Eq.
+14, using DESI DR2 data with the FRB–based $H(z)$ and propagated BAO uncertainties.
+Colored curves show the FRB–only estimator $D_{A}^{\rm FRB}(z;\Omega_{k})$ evaluated at the mean $\Omega_{k}$ of each analysis; the shaded bands denote credible regions.
+From the covariance-based analysis by employing our 120 FRBs and the DESI DR2 datasets, we derive an estimated $\Omega_{k}=-0.31\pm 0.57$ for the cosmic curvature parameter.
+In contrast, neglecting the off-diagonal elements of the covariance matrix yields a bit tighter posterior distribution, $\Omega_{k}=-0.13\pm 0.46$.
 Despite variations in central values, all estimates remain consistent with a spatially flat Universe ($\Omega_{k}=0$) at the $1\sigma$ level.
 The integration of FRB-based comoving distance reconstruction serves as an independent and complementary cosmological probe, reinforcing the growing utility of FRBs in precision cosmology.
+In panel (b) of Figure 3 we compare the FRB-only reconstruction of the angular–diameter distance, $D_{A}^{\rm FRB}(z;\Omega_{k})$, obtained with our two likelihood treatments against the discrete $D_{A}^{\rm BAO+FRB}$ estimates at the DESI DR2 effective redshifts – black squares.
+Solid curves show the posterior–mean $D_{A}^{\rm FRB}(z)$ for each treatment, while the shaded bands indicate the propagated $3\sigma$ uncertainty, which naturally widens toward high $z$ where the FRB sample is sparse.
+The DESI points (with $1\sigma$ error bars) are derived from the measured BAO ratios via Eq.
+14 combined with the FRB-based $H(z)$ reconstruction.
+The two FRB-only solutions agree closely over $0.4\!\lesssim\!z\!\lesssim\!2$, with the COV band being broader—as expected when correlations are retained—yet both envelopes encompass the BAO+FRB determinations within uncertainties.
+The peak around $z\!\sim\!1.6$ is consistent with the expected maximum of $D_{A}(z)$ in late-time cosmologies, and the overall consistency of the three estimators supports our curvature analysis.
 Taken together, these results demonstrate that ANN reconstructions of comoving distances from FRBs, when combined with state-of-the-art BAO measurements, provide competitive constraints on spatial curvature, comparable to those obtained from other late-time cosmological probes Liu et al.
@@
 (2021).
-While Gaussian-based methods yield tighter intervals, they may underestimate uncertainties due to neglected correlations.
-The apparent preference for negative values in this case may result from omitting the covariance matrix, which can bias the final result.
-In contrast, the approach that incorporates the full covariance matrix, though yielding broader confidence regions, provides a more conservative and robust assessment of parameter uncertainties.
-These findings highlight the critical role of accounting for the full covariance structure in cosmological parameter inference, especially when combining different datasets or when intrinsic correlations are present.
+In our analysis, both likelihood treatments are statistically consistent with spatial flatness at the $1\sigma$ level, but they behave differently with respect to uncertainty budgeting.
+The Gaussian approximation produces visibly tighter credible intervals, yet it risks understating the true errors because it neglects cross–terms induced by the common FRB–based $H(z)$ entering both $D_{A}$ and $D_{C}$.
+When we retain the full covariance, the posteriors broaden as expected, yielding a more conservative and robust assessment of parameter uncertainties.
+Importantly, the mild preference for $\Omega_{k}<0$ persists in both treatments, and the shift of the posterior mean when toggling the covariance is small, indicating that our qualitative conclusion is not driven by the covariance choice.
+Figure 4: Constraints on the spatial curvature $\Omega_{k}$.
+Horizontal bars show $1\sigma$ confidence intervals for our FRB + BAO (COV) and FRB + BAO (GAUSS) analyses, alongside a set of literature determinations (Planck 2018 + BAO; Planck 2018 (no lensing); EFTofLSS/BAO; BAO + Pantheon+; CC + SGL; Planck 2018 + $C$; etc.).
+The vertical dashed line marks $\Omega_{k}=0$.
+Incorporating the full covariance (COV) yields broader, more conservative constraints than the Gaussian approximation (GAUSS), yet both remain consistent with spatial flatness at $1\sigma$.
 […]
@@
 This dual-path approach enabled robust constraints on the curvature parameter $\Omega_{k}$, independent of early-universe assumptions or fiducial cosmological models.
+Our analysis yields $\Omega_{k}=-0.31\pm 0.57$ (full covariance) and $\Omega_{k}=-0.13\pm 0.46$ (Gaussian approximation).
+While the uncorrelated treatment produces tighter intervals, it risks underestimating uncertainties, we therefore adopt the covariance-based constraint as our headline result.
 Our results indicate that, within current observational limits, all estimates remain consistent with a spatially flat Universe.
@@
 Acknowledgements.
-JASF thanks FAPES for their financial support.
+JASF thanks FAPES and the National Research Foundation (NRF) for their financial support.
 WSHR acknowledges partial financial support from FAPES.
@@
 acknowledges financial support from the State Agency for Research of the Spanish Ministry of Science and Innovation under grant PID2022-136828NB-C41/AEI/10.13039/501100011033/, and by “ERDF A way of making Europe”, by the “European Union”.
-He also thanks support from PIP 0554 (CONICET).
+He also thanks the support from PIP 0554 (CONICET).
+JASF thanks Amanda Weltman and Surajit Kalita for fruitful discussions.
+.1 Dataset
+In this appendix we present the Full Localized FRBs sample, including the ones used in this work and those we neglected for the reasons described in the main text.
+Reported observables are the measured dispersion measure $\mathrm{DM}_{\rm obs}$ and the Milky Way contribution from the NE2001 model.
+Table 2: Localized FRBs
+FRB $z$ $\mathrm{DM}_{\rm obs}$ $\mathrm{DM}_{\rm MW}$ (NE2001) Ref.
+FRB200120E 0.0008 87.8 40.8 Connor et al. (2024)
+FRB181030A 0.00385 103.396 41.1 Lanman et al. (2022)
+FRB171020A 0.008672 114.1 38.0 Mahony et al. (2018)
+FRB220319D 0.011228 110.98 65.25 Ravi et al. (2023)
+FRB231129A 0.019000 198.50 58.3 Amiri et al. (2025)
+FRB240210A 0.023686 283.73 28.7 Shannon et al. (2025)
+FRB181220A 0.027460 208.66 118.5 Bhardwaj et al. (2024)
+FRB231230A 0.029800 131.40 23.0 Amiri et al. (2025)
+FRB181223C 0.030240 111.61 19.9 Bhardwaj et al. (2024)
+FRB190425A 0.03122 127.78 48.7 Bhardwaj et al. (2024)
+FRB180916B 0.033700 349.35 199.0 Marcote et al. (2020)
+FRB230718A 0.0357 477.0 420.6 Glowacki et al. (2024)
+FRB231120A 0.036800 437.74 43.8 Sharma et al. (2024)
+FRB240201A 0.042729 374.50 38.6 Shannon et al. (2025)
+FRB220207C 0.043040 262.38 76.1 Law et al. (2024)
+FRB211127I 0.046900 234.83 42.5 Glowacki et al. (2023)
+FRB201123A 0.050700 433.55 251.7 Rajwade et al. (2022)
+FRB230926A 0.055300 222.80 52.6 Amiri et al. (2025)
+FRB200223B 0.060240 201.80 45.6 Ibik et al. (2024)
+FRB190303A 0.064000 223.20 29.8 Michilli et al. (2023)
+FRB231204A 0.064400 221.00 34.9 Amiri et al. (2025)
+FRB231206A 0.065900 457.70 59.1 Amiri et al. (2025)
+FRB210405I 0.066 565.17 516.1 Driessen et al. (2024)
+FRB180814A 0.068 190.9 87.5 Michilli et al. (2023)
+FRB211212A 0.070700 206.00 38.8 et al.
+FRB231005A 0.071300 189.40 33.5 Amiri et al. (2025)
+FRB190418A 0.071320 182.78 70.2 Bhardwaj et al. (2024)
+FRB231123A 0.072900 302.10 89.7 Amiri et al. (2025)
+FRB220912A 0.0771 219.46 125.2 Ravi et al. (2023)
+FRB231011A 0.078300 186.30 70.4 Amiri et al. (2025)
+FRB220509G 0.089400 269.53 55.6 Law et al. (2024)
+FRB230124A 0.093900 590.57 38.6 Sharma et al. (2024)
+FRB201124A 0.098000 413.52 139.9 Ravi et al. (2022)
+FRB230708A 0.105000 411.51 60.3 Shannon et al. (2025)
+FRB231223C 0.105900 165.80 47.9 Amiri et al. (2025)
+FRB191106C 0.107750 332.20 25.0 Ibik et al. (2024)
+FRB231128A 0.107900 331.60 64.7 Amiri et al. (2025)
+FRB230222B 0.110000 187.80 27.7 Amiri et al. (2025)
+FRB231201A 0.111900 169.40 70.0 Amiri et al. (2025)
+FRB220914A 0.113900 631.28 54.7 Law et al. (2024)
+FRB190608B 0.117780 338.70 37.3 Chittidi et al. (2021)
+FRB230703A 0.118400 291.30 26.9 Amiri et al. (2025)
+FRB240213A 0.118500 357.40 40.0 Connor et al. (2024)
+FRB230222A 0.122300 706.10 134.2 Amiri et al. (2025)
+FRB190110C 0.122440 221.60 37.1 Ibik et al. (2024)
+FRB230628A 0.127000 344.95 39.0 Sharma et al. (2024)
+FRB240310A 0.127000 601.80 30.1 Shannon et al. (2025)
+FRB210807D 0.129300 251.90 121.2 Glowacki et al. (2023)
+FRB240114A 0.130000 527.65 49.7 Kumar et al. (2024)
+FRB240209A 0.138400 176.57 55.5 Shah et al. (2025)
+FRB210410D 0.141500 571.20 56.2 Caleb et al. (2023)
+FRB230203A 0.146400 420.10 67.3 Amiri et al. (2025)
+FRB231226A 0.156900 329.90 38.1 Shannon et al. (2025)
+FRB230526A 0.157000 361.40 31.9 Shannon et al. (2025)
+FRB220920A 0.158239 314.99 39.9 Law et al. (2024)
+FRB200430A 0.160800 380.10 27.2 Heintz et al. (2020)
+FRB190203A 0.170000 134.40 20.3 Tyul’bashev et al. (2025)
+FRB210603A 0.177200 500.15 39.5 Cassanelli et al. (2024)
+FRB220529A 0.183900 246.00 40.0 Li et al. (2025)
+FRB230311A 0.191800 364.30 92.5 Amiri et al. (2025)
+FRB220725A 0.192600 290.40 30.7 Shannon et al. (2025)
+FRB121102A 0.192730 557.00 188.4 Tendulkar et al. (2017)
+FRB221106A 0.204400 343.80 34.8 Shannon et al. (2025)
+FRB240215A 0.210000 549.50 47.9 Connor et al. (2024)
+FRB230730A 0.211500 312.50 85.2 Amiri et al. (2025)
+FRB210117A 0.214500 729.10 34.4 Bhandari et al. (2023)
+FRB221027A 0.229000 452.50 47.2 Sharma et al. (2024)
+FRB191001A 0.234000 506.92 44.2 Heintz et al. (2020)
+FRB190714A 0.236500 504.13 38.5 Heintz et al. (2020)
+FRB221101B 0.239500 491.55 131.2 Sharma et al. (2024)
+FRB220825A 0.241397 651.24 78.5 Law et al. (2024)
+FRB190520B 0.2418 1204.7 63 Ocker et al. (2022)
+FRB191228A 0.243200 297.50 32.9 Bhandari et al. (2022)
+FRB231017A 0.245000 344.20 31.2 Amiri et al. (2025)
+FRB220307B 0.248123 499.27 128.2 Law et al. (2024)
+FRB221113A 0.250500 411.03 91.7 Sharma et al. (2024)
+FRB220831A 0.262 1146.25 126.8 Connor et al. (2024)
+FRB231123B 0.262100 396.86 40.3 Amiri et al. (2025)
+FRB230307A 0.270600 608.85 37.6 Sharma et al. (2024)
+FRB221116A 0.276400 643.45 132.3 Sharma et al. (2024)
+FRB220105A 0.278500 583.00 22.0 Shannon et al. (2025)
+FRB210320C 0.279700 384.80 39.3 Glowacki et al. (2023)
+FRB221012A 0.284669 441.08 54.3 Law et al. (2024)
+FRB240229A 0.287000 491.15 38.0 Connor et al. (2024)
+FRB190102C 0.291200 364.50 57.4 Macquart et al. (2020)
+FRB220506D 0.300390 396.97 84.6 Law et al. (2024)
+FRB230501A 0.301500 532.47 125.7 Sharma et al. (2024)
+FRB230503E 0.320000 483.74 88.0 Pastor-Marazuela et al. (2025)
+FRB180924B 0.321200 361.42 40.5 Bannister et al. (2019)
+FRB231025B 0.323800 368.70 48.6 Amiri et al. (2025)
+FRB230125D 0.3265 640.08 88.0 Pastor-Marazuela et al. (2025)
+FRB230626A 0.327 452.723 39.3 Sharma et al. (2024)
+FRB180301A 0.330400 552.00 151.7 Price et al. (2019)
+FRB231220A 0.335500 491.20 49.9 Connor et al. (2024)
+FRB211203C 0.343900 636.20 63.7 Shannon et al. (2025)
+FRB230808F 0.347200 653.20 27.0 Hanmer et al. (2025)
+FRB220208A 0.351000 437.00 101.6 Sharma et al. (2024)
+FRB230902A 0.361900 440.10 34.1 Shannon et al. (2025)
+FRB220726A 0.361900 686.23 89.5 Sharma et al. (2024)
+FRB220717A 0.362950 637.34 118.3 Rajwade et al. (2024)
+FRB200906A 0.3688 577.8 35.8 Bhandari et al. (2022)
+FRB220330D 0.3714 468.1 38.6 Sharma et al. (2024)
+FRB240119A 0.376000 483.10 38.0 Connor et al. (2024)
+FRB190611B 0.377800 321.40 57.8 Heintz et al. (2020)
+FRB220501C 0.381000 449.50 30.6 Shannon et al. (2025)
+FRB230613A 0.392300 483.51 30.0 Pastor-Marazuela et al. (2025)
+FRB220204A 0.401200 612.58 50.7 Connor et al. (2025)
+FRB230712A 0.452500 587.57 39.2 Sharma et al. (2024)
+FRB230907D 0.463800 1030.79 29.0 Pastor-Marazuela et al. (2025)
+FRB181112A 0.475500 589.27 102.0 Prochaska et al. (2019)
+FRB231020B 0.477500 952.20 34.0 Pastor-Marazuela et al. (2025)
+FRB220307B 0.248123 499.27 128.2 Law et al. (2024)
+FRB220310F 0.477958 462.24 46.3 Law et al. (2024)
+FRB220918A 0.491000 656.80 41.0 Shannon et al. (2025)
+FRB231210F 0.500000 720.60 32.0 Pastor-Marazuela et al. (2025)
+FRB190711A 0.522000 593.10 56.5 Heintz et al. (2020)
+FRB230216A 0.531000 828.00 38.5 Sharma et al. (2024)
+FRB221219A 0.553000 706.71 44.4 Sharma et al. (2024)
+FRB230814A 0.553000 696.40 104.8 Connor et al. (2024)
+FRB190614D 0.600000 959.20 87.8 Law et al. (2024)
+FRB210924D 0.600000 737.00 45.0 Carli et al. (2024)
+FRB231010A 0.610000 442.59 41.0 Pastor-Marazuela et al. (2025)
+FRB220418A 0.622000 623.25 36.7 Law et al. (2024)
+FRB220224C 0.627100 1140.20 52.0 Sharma et al. (2024)
+FRB190523A 0.660000 760.80 37.2 Ravi et al. (2019)
+FRB190208A 0.830000 580.01 71.5 Hewitt et al. (2024)
+FRB220222C 0.853000 1071.20 56.0 Pastor-Marazuela et al. (2025)
+FRB240123A 0.968000 1462.00 90.2 Connor et al. (2024)
+FRB221029A 0.975000 1391.75 43.8 Sharma et al. (2024)
+FRB220610A 1.016000 1458.15 31.0 Gordon et al. (2024)
+FRB230521B 1.354000 1342.90 138.8 Connor et al. (2024)
+FRB240304B 2.148000 2330.00 28.1 Caleb et al. (2025)